# Retrieval-Based Video Language Model for Efficient Long Video Question Answering

### v1

@@ Abstract
The remarkable natural language understanding, reasoning, and generation capabilities of large language models (LLMs) have made them attractive for application to video question answering (Video QA) tasks, utilizing video tokens as contextual input. However, employing LLMs for long video understanding presents significant challenges and remains under-explored. The extensive number of video tokens leads to considerable computational costs for LLMs while using aggregated tokens results in loss of vision details. Moreover, the presence of abundant question-irrelevant tokens introduces noise to the video QA process. To address these issues, we introduce a simple yet effective retrieval-based video language model (R-VLM) for efficient and interpretable long video QA. Specifically, given a question (query) and a long video, our model identifies and selects the most relevant $K$ video chunks and uses their associated visual tokens to serve as context for the LLM inference. This effectively reduces the number of video tokens, eliminates noise interference, and enhances system performance. Our experimental results validate the effectiveness of our framework for comprehending long videos. Furthermore, based on the retrieved chunks, our model is interpretable that provides the justifications on where we get the answers.
1 Introduction
@@
There is a growing trend towards leveraging large language models (LLMs) for video QA (Maaz et al., 2023; Wang et al., 2023; Li et al., 2023b; Zhang et al., 2023b). On one hand, LLMs benefit from the vast knowledge acquired through training on enormous text corpora; on the other hand, they provide users with a more natural and intuitive way to interact with video data. Generally, vision tokens extracted from a video snippet are transformed and used as input (prompt) to the LLM, along with the text query. However, it is worth noting that the consumption of abundant vision tokens by a LLM can significantly increase the memory and computational burden, making it unaffordable for low-resource GPU agents. To mitigate this issue, Maaz et al. (2023) perform global spatial and temporal pooling on the video tokens, although this comes at the cost of losing detail due to pooling. Zhang et al. (2023b) aggregate video tokens using a Q-former with cross attention. Most of these methods are designed primarily for short-video QA tasks, where answer-related frames usually spread over the trimmed video snippet.
In practical application scenarios, users often pose flexible questions related to long videos (e.g., longer than 1 minute), in which the segments containing pertinent information for answering the questions constitute merely a small fraction of the entire video. The presence of answer-irrelevant tokens are redundancy and may potentially interfere with the video QA process, diminishing its effectiveness. Therefore, it is imperative to develop a simple yet efficient framework that can handle long video QA tasks.
To address these challenges, we draw inspiration from biology and cognitive science. As we know, human working memory is a cognitive system responsible for temporarily holding and manipulating information necessary for complex tasks such as reasoning, learning, and comprehension (Thompson & Madigan, 2013). Faced with the vast amount of information stored in long-term memory, working memory selectively retrieves relevant information while filtering out irrelevant data for further cognition. Motivated by this, we aim to design a framework that is capable of identifying and concentrating on relevant video segments while filtering out irrelevant information, ensuring accurate and efficient question answering without imposing excessive computational demands.
@@
Figure 1: Illustration of our proposed retrieval-based video language model for efficient long video question answering. We encode an input long video into a sequence of video chunks, with each chunk represented by a set of spatial and temporal visual tokens. Question-guided retrieval is performed to find the top $K$ relevant video chunks, with their tokens as the context/prompt of the LLM for answer generation. Soft matching (SM) loss is used to regularize the retrieval related learning.
@@
In this paper, we propose a retrieval-based video-language model, R-VLM, for efficient and interpretable long-video question answering. Fig. 1 shows the overall framework. Specifically, given a long video and a question (query), we divide the video into a sequence of non-overlapping video chunks, with each chunk representing a short video segment, e.g., 4 seconds with a sample rate of 1fps. Note that we sample at a low frame rate in considering the memory limitation and video redundancy. To allow a chunk to contain dynamic temporal information, we use 4 seconds that are expected to contain temporal dynamics as the chunk unit. We then aggregate the encoded tokens within a chunk through spatial and temporal pooling to obtain chunk tokens, reducing some redundancy while preserving considerable local details. We perform question-guided retrieval to obtain the top $K$ most relevant chunks, and subsequently use the small set of tokens (after projection) from these chunks as input for the frozen LLM for video QA inference. In this way, we efficiently match and pass the most question-pertinent visual information to the LLM for inference.
Our framework demonstrates superior zero-shot generalization performance on several video QA datasets, outperforming the baseline method Video-ChatGPT (Maaz et al., 2023) by 6.8%, 2.8%, 4.8%, 6.0% in accuracy on the WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0 dataset, respectively. Notably, the retrieved video chunks provide justification for the model’s responses, offering interpretability to the video language model on where matters for answering the question.
@@
  • We propose a retrieval-based video language model for efficient long video understanding. To the best of our knowledge, we are the first time to validate the feasibility of using retrieval for long video question answering with large video-language model.
  • Thanks to the retrieval-based context/prompt chunk selection mechanism, our model is more interpretable, where the selected chunks provide insight on based on where and why the model generates the answer.
  • Our method achieves significant performance improvement over the baseline method. The ablation studies demonstrate the superiority of our retrieval mechanism.
2 Related Work
Large Language Models. LLMs (Radford et al., 2018; Brown et al., 2020; Ouyang et al., 2022) have experienced significant advances and demonstrated remarkable capabilities such as language understanding, reasoning, generation, and in context learning. These abilities enable LLMs to tackle diverse tasks with user prompts in a zero-shot fashion, reflecting their powerful adaptability and generalization. The Instruction tuning strategy as used in ChatGPT (OpenAI, 2023) improves the model’s alignment with user intentions and optimizes the quality of generation. At the same time, many open-source instruction tuned models have emerged, such as LLaMA (Touvron et al., 2023), OPT (Zhang et al., 2022) and GML (Zeng et al., 2022), which have greatly promoted technological advancement and made significant contributions to the community.
Vision-Language Models with Pre-trained LLMs. Recent advances in LLMs have inspired and promoted the integration of pre-trained LLMs with visual processing components for multimodal understanding (Alayrac et al., 2022; Li et al., 2023a; Zhu et al., 2023; Liu et al., 2023; Huang et al., 2023). Flamingo (Alayrac et al., 2022) and BLIP-2 (Li et al., 2023a) are seminal works that leverage the pre-trained LLM and CLIP image encoder for zero-shot image-text abilities. BLIP2 (Li et al., 2023a) introduces a Q-Former to map the image tokens from the CLIP encoder (Radford et al., 2021) to the textual embedding space of LLMs. LLaVA (Liu et al., 2023) maps the image spatial tokens to the textual embedding space of LLMs using a linear projector. LLaVA (Liu et al., 2023), MiniGPT4 (Zhu et al., 2023), and mPLUG-owl (Ye et al., 2023) promote instruction following using image-instruction-following datasets.
Some works have focused on enabling video language understanding by integrating vision encoder and LLMs (Zhang et al., 2023a; Maaz et al., 2023; Li et al., 2023b; Wang et al., 2023). Video-LLaMA (Zhang et al., 2023a) uses a video Q-former to assemble the features from the pre-trained CLIP image encoder. Video-ChatGPT (Maaz et al., 2023) performs spatial and temporal pooling of the feature maps encoded by the pre-trained CLIP image encoder. The ensembled vision tokens are taken as input to the LLMs. Most of these works globally aggregate video tokens. They may work well for short videos. But for the long video language comprehension, they would be less efficient. The information of the question related video segments may be submerged to the global-wise token representations, making the reasoning difficult. We aim to design a simple yet efficient video language model for long video understanding. Chunk-wise representations and learnable retrieval are introduce to address the challenges.
@@ 2 Related Work
3 Proposed Retrieval-based Video Language Model
@@
Given a lengthy video, it is time-consuming to watch the entire content to obtain the desired information. A Video QA system, which can automatically infer answers based on the long video and user’s query (question), is in high demand. LLMs possess extensive world knowledge and reasoning capabilities, albeit at the expense of high computational complexity. Some previous works on multi-modal language models utilize a set of projected vision tokens as input (context) to LLM for inference (Li et al., 2023a; Zhu et al., 2023), where the inference cost is proportional to the number of input vision tokens.
@@
Leveraging powerful LLMs to understand long videos presents challenges. Firstly, we expect to use only a small set of tokens as input to reduce computational costs. Secondly, as the video length increases, in general, there is a corresponding growth in the number of vision tokens and the overall amount of information. Representing a long video with very few tokens becomes difficult. Furthermore, question-irrelevant information may interfere with answer generation. Motivated by the retrieval mechanism of brain, we introduce a question-guided retrieval mechanism to identify and select a few relevant video chunks as context of LLM.
@@
Fig. 1 illustrates the overall framework of our proposed retrieval-based video language model (R -VLM), designed for efficient long video question answering. The framework comprises several components: a frozen LLM, a frozen image encoder, a frozen text encoder, a memory for storing video chunk tokens, a retrieval module for selecting the question-relevant chunks, a learnable MLP block, and a projector. Through end-to-end training with cross-entropy loss and the proposed soft matching (SM) loss, the MLP block is optimized to learn to identify the most relevant $K$ chunks, while the projector is optimized to align the selected vision tokens with the text (question) space. In the following subsections, we provide a detailed explanation of the key components and designs.
3.1 Video Tokenization
Video tokenization involves encoding the raw video into vision tokens, which will be processed and then passed to the LLM for inference.
@@ 3.1 Video Tokenization
Vision token extraction of a chunk. The vision tokens of a chunk are obtained by encoding the images and performing spatial-temporal pooling. Following Alayrac et al. (2022) and Maaz et al. (2023), we adapt the pretrained language-image model CLIP (Radford et al., 2021) to extract per frame vision token features, which are suited for capturing high-level semantic information. For the $j^{th}$ video chunk $V_{i}^{j}\in\mathbb{R}^{M\times H\times W\times C}$, we obtain $M\times h\times w$ vision tokens extracted by the CLIP vision encoder, where $h=H/p,w=W/p$. $p$ denotes the patch size (which is 14 for ViT-L/14).
The preliminary token number of a chunk is large, i.e., $M\times h\times w=4\times 16\times 16=1024$. This would result in a high computational burden and large memory requirement for the LLM even though we only select a few chunks as input to LLM. We found that reducing the spatial resolution by a factor of 4 leads to marginal difference in performance while this can significantly reduce the number of tokens by 75%. Therefore, we perform spatial average pooling with stride 2 to have $M$$\times$ $\bar{h}$$\times$ $\bar{w}$ = $4\times 8\times 8=256$ tokens per chunk, where $\bar{h}$=$h/2$ and $\bar{w}$=$w/2$. This is equivalent to that we take the CLIP features of reduced resolution as the extracted feature. How to further reduce the number of tokens while preserving spatial-temporal features of a chunk? Motivated by Maaz et al. (2023), we perform spatial-temporal pooling on the token tensor. Particularly, global spatial average pooling for each frame is performed and thus we obtain $M$ tokens for the $M$ frames (e.g., 4 tokens). Temporal pooling for each spatial position is performed to have $N$= $\bar{h}$$\times$ $\bar{w}$ = 8 $\times$ 8 = 64 tokens. We have $N+M=64+4=68$ tokens $F_{i}^{j}\in\mathbb{R}^{(N+M)\times D}$, with $D$ dimension for each token. Compared with the original 1024 in a chunk, the number of tokens has been reduced to 6.6%, which brings high calculation efficiency for the later LLM inference.
In contrast to Maaz et al. (2023), which performs global spatial and temporal pooling over the entire video, ours with chunks is capable of preserving local details and is better suited for long video QA.
3.2 Question-guided Retrieval for Chunk Selection
@@
where the MLP block is a two-layer perceptron consisting of two fully connected layers and an ReLU activation function in between. This MLP transformation strengthens the correlation between the question representation and the potential corresponding chunk features for chunk selection. Achieving this is non-trivial, as we do not have any ground-truth locations of the question relevant chunks for supervision.
To identify the question-related chunks, we measure the affinity between the text feature vector and the chunk representation. For simplicity, we obtain the representation feature of the $j^{th}$ chunk for the matching by aggregating its $N+M$ vision tokens by averaging pooling as $\mathbf{v}_{i}^{j}=Avgpool(F_{i}^{j})$. This is parameter-free, avoiding the need for large memory in optimization. We store the chunk representation features and chunk tokens in a memory to facilitate retrieval.
We compute the similarity scores between the question representation $\mathbf{q}_{i}$ and each video chunk representation $\mathbf{v}_{i}^{j}$ using cosine similarity metric as
$s_{i}^{j}=\cos(\mathbf{q}_{i},\mathbf{v}_{i}^{j})=\frac{\mathbf{q}_{i}\cdot\mathbf{v}_{i}^{j}}{\lVert\mathbf{q}_{i}\rVert\lVert\mathbf{v}_{i}^{j}\rVert}.$ (2)
@@
We rank the video chunks based on their similarity scores $s_{i}^{j}$ (where $j=1,\cdots,L_{i}$) and select the top $K$ most relevant chunks. The $K\times(N+M)$ vision tokens of these chunks are input to the LLM after a linear projection (a fully connected layer) on each token.
3.3 End-to-End Optimization

### v2

# Long Video Understanding with Learnable Retrieval in Video-Language Models

@@ Abstract
The remarkable natural language understanding, reasoning, and generation capabilities of large language models (LLMs) have made them attractive for application to video understanding, utilizing video tokens as contextual input. However, employing LLMs for long video understanding presents significant challenges. The extensive number of video tokens leads to considerable computational costs for LLMs while using aggregated tokens results in loss of vision details. Moreover, the presence of abundant question-irrelevant tokens introduces noise to the video reasoning process. To address these issues, we introduce a simple yet effective learnable retrieval-based video-language model (R-VLM) for efficient long video understanding. Specifically, given a question (query) and a long video, our model identifies and selects the most relevant $K$ video chunks and uses their associated visual tokens to serve as context for the LLM inference. This effectively reduces the number of video tokens, eliminates noise interference, and enhances system performance. We achieve this by incorporating a learnable lightweight MLP block to facilitate the efficient retrieval of question-relevant chunks, through the end-to-end training of our video-language model with a proposed soft matching loss. Our experimental results on multiple zero-shot video question answering datasets validate the effectiveness of our framework for comprehending long videos.
Index Terms:
Large Language Models, Long Video Question Answering, Retrieval-based Video-language Model.
I Introduction
@@
There is a growing trend towards leveraging large language models (LLMs) for video QA [1, 2, 3, 4]. On one hand, LLMs benefit from the vast knowledge acquired through training on enormous text corpora; on the other hand, they provide users with a more natural and intuitive way to interact with video data. Generally, visual tokens extracted from a video snippet are transformed and used as input (prompt) to the LLM, along with the text query. However, it is worth noting that the consumption of abundant visual tokens by a LLM can significantly increase the memory and computational burden, making it unaffordable for low-resource GPU agents. To mitigate this issue, Video-ChatGPT [1] performs global spatial and temporal pooling on the video tokens, although this comes at the cost of losing detail due to pooling. Video-LLaMA [4] aggregates video tokens using a Q-former with cross attention. Most of these methods are designed for short-video QA tasks, where answer-related frames usually spread over the trimmed video snippet.
@@
Figure 1: Illustration of our learnable retrieval-based video-language model for efficient long video question answering. We encode an input long video into a sequence of video chunks, with each chunk represented by a set of spatial and temporal visual tokens. Question-guided retrieval is performed to find the top $K$ relevant video chunks, with their tokens as the input to the LLM for answer generation. Here, we use $K=1$ for illustration purpose. A learnable lightweight MLP block (following the text encoder) and the projector are trained end-to-end, where the encoders and LLM are frozen. Soft matching (SM) loss is introduced to regularize the retrieval related learning.
In practical application scenarios, users often raise questions about long videos (e.g., longer than 1 minute), in which the segments that are relevant to the questions usually constitute only a small fraction of the entire video. The presence of answer-irrelevant segments constitutes redundancy and can potentially interfere with the video QA process, thus reducing its effectiveness. Therefore, it is imperative to develop a simple yet efficient framework for long video QA.
To address these challenges, we draw inspiration from biology and cognitive science. As we know, human working memory is a cognitive system responsible for temporarily holding and manipulating information necessary for complex tasks such as reasoning, learning, and comprehension [5]. Faced with the vast amount of information stored in long-term memory, working memory selectively retrieves relevant information while filtering out irrelevant data for further cognition. Motivated by this, we aim to design a framework that is capable of identifying and concentrating on relevant video segments while filtering out irrelevant information, ensuring accurate and efficient question answering without imposing excessive computational demands.
@@
In this paper, we propose a retrieval-based video-language model, R-VLM, for effective long-video question answering. Fig. 1 shows the overall framework. Specifically, given a long video and a question (query), we divide the video into a sequence of non-overlapping video chunks, with each chunk representing a short video segment (e.g., 4 seconds). Note that we sample at a low frame rate in considering the memory limitation and video redundancy. To allow a chunk to contain dynamic temporal information, we use 4 seconds that are expected to contain temporal dynamics as the chunk unit. We then aggregate the encoded tokens within a chunk through spatial and temporal pooling to obtain chunk tokens, reducing redundancy while preserving considerable local details. We perform question-guided retrieval by introducing a lightweight learnable MLP block on top of a CLIP text encoder to obtain the top $K$ most relevant chunks, and subsequently use the small set of tokens (after projection) from these chunks as input to the LLM for video reasoning. In this way, we efficiently match and pass the most question-pertinent visual information to the LLM for inference.
Our framework demonstrates superior zero-shot generalization performance on multiple video QA datasets,
@@
  • We propose a learnable retrieval-based video-language model for effective long video understanding.
  • We validate the feasibility of using learnable lightweight retrieval for selecting the most question-relevant video chunks to answer questions from a long video, within large video-language model in an end-to-end manner.
  • Our method achieves superior performance and the ablation studies demonstrate the superiority of our retrieval mechanism.
II Related Work
Video Question Answering (Video QA) has emerged as a crucial task for extracting information from videos through natural language interactions. Before the prevalence of LLM, conventional methods focus on network and multi-modal interaction designs for video understanding [6, 7, 8, 9, 10, 11, 12, 13]. Yuan et al.[6] propose AMN to learn multi-modal feature representations by finding a more coherent subspace for video clips and the corresponding texts based on GANs. Wang et al.[7] propose an alternating attention network, which iteratively focuses on frame regions, video frames, and words in the question in multi-turns to generate better joint representations of the video and question. DualVGR [8] uses a multi-view graph attention network to provide a context-aware feature representation. A query punishment module is applied to modulate irrelevant visual features through multiple reasoning cycles to reduce their interference. Qian et al.[9] propose a scheme to localize the question to a segment in the video to infer the answer, where a 3DCNN network and attention modules are designed to extract features and generate location proposals. Based on existing LLM and visual encoder, how to design a simple framework that can leverage the LLM and visual encoder for long video understanding is still under-explored. We propose R-VLM that leverage a lightweight MLP block to enable a learnable retrieval module for efficient video question answering.
Recent advances in LLMs have inspired and promoted the integration of pre-trained LLMs with visual processing components for multimodal understanding [14], [15], [16], [17], [18], [19]. Flamingo [14] and BLIP-2 [15] are seminal works that leverage the pre-trained LLM and CLIP image encoder for zero-shot image-text abilities. BLIP2 [15] introduces a Q-Former to map the image tokens from the CLIP encoder [20] to the textual embedding space of LLMs. InstructBLIP [21] uses instruction-aware Q-Former to globally aggregate the visual tokens. LLaVA [17] maps the image spatial tokens to the textual embedding space of LLMs using a linear projector. LLaVA [17], MiniGPT4 [16], and mPLUG-owl [22] promote instruction following using image-instruction-following datasets.
Some works have focused on enabling video-language understanding by integrating vision encoder and LLMs [23], [1], [3], [2]. Video-LLaMA [23] uses a video Q-former to assemble the features from the pre-trained CLIP image encoder. Video-ChatGPT [1] performs spatial and temporal pooling of the feature maps encoded by the pre-trained CLIP image encoder. The ensembled visual tokens are taken as input to the LLMs. Most of these works globally aggregate video tokens. They may work well for short videos. But for the long video-language comprehension, they would be less efficient. The information of the question-related video segments may be submerged to the global-wise token representations, making the reasoning difficult.
SeViLA [24] utilizes the BLIP-2 image-language model for selecting temporal key frames which are subsequently inputted into BLIP-2 for video-based question answering. The process involves complex training for convergence and the frame selection is computationally expensive due to the necessity of processing potential frames through the large BLIP-2 model (with even the smallest version being 3.1 billion parameters). This would not be suitable for long video understanding.
In order to understand long videos, MovieChat [25] leverages short-term and long-term memories for video representation. LLaMA-VID [26] represents every frame coarsely by two tokens to reduce the computation and memory cost for long video reasoning, which however still introduces temporal redundancy and interference in LLM inference. SeViT [27] proposes a framework for long untrimmed video understanding. It includes a non-parametric frame retriever to select a few query-relevant frames from the video and uses them as the input to a video-grounded text generator. However, the frame selector is non-parametric and the system cannot be end-to-end optimized, resulting in the inferiority of the selected frames and sub-optimal performance. Different from SeViT, we introduce learnable retrieval through a light-weighted MLP block for building an end-to-end video-language model, powered by LLM. In [28], a Frame-Prompter is designed to select key video frames. However, with only video frames as input to the Frame-Prompter, the selection is not question aware and would lead to ineffective. Wang et al.[29] fuse the question-and-answer pairs as event descriptions to find keyframes as target moments and pseudo-labels. Unlike [29], we do not use pseudo-labels for supervision. Our framework leverages the end-to-end learning to automatically select the question-relevant chunks, providing a large optimization space and avoiding the side effects from the inaccuracy of pseudo-labels. VideoTree [30] extracts query-relevant information from the input video through an iterative process, based on their relevance to the query scored by an LLM. However, leveraging an additional LLM for key frame selection increases the complexity of the system. In contrast, our retrieval design is lightweighted and end-to-end trained in the VLM.
III Retrieval-based Video-Language Model
@@
Given a lengthy video, it is time-consuming to watch the entire content to obtain the desired information. A video understanding system, which can automatically infer answers based on the long video and user’s query (question), is in high demand. LLMs possess extensive world knowledge and reasoning capabilities, albeit at the expense of high computational complexity. Some previous works on multi-modal language models utilize a set of projected visual tokens as input (context) to LLM for inference [15], [16], where the inference cost is proportional to the number of input visual tokens.
@@
Leveraging powerful LLMs to understand long videos presents challenges. Firstly, we expect to use only a small set of tokens as input to reduce computational costs. Secondly, as the video length increases, in general, there is a corresponding growth in the number of visual tokens and the overall amount of information. Representing a long video with very few tokens becomes difficult. Furthermore, question-irrelevant information may interfere with reasoning. Motivated by the retrieval mechanism of brain, we introduce a question-guided retrieval mechanism to identify and select a few relevant video chunks as context of LLM.
@@
Fig. 1 illustrates the overall framework of our proposed retrieval-based video-language model (R-VLM), designed for efficient long video understanding. The framework is comprised by several components: a frozen LLM, a frozen image encoder, a frozen text encoder, a memory for storing video chunk tokens, a retrieval module for selecting the question-relevant chunks, a learnable MLP block, and a projector. Through end-to-end training with cross-entropy loss and the proposed soft matching (SM) loss, the MLP block is optimized to learn to identify the most relevant $K$ chunks, while the projector is optimized to align the selected visual tokens with the text (question) space. In the following subsections, we provide a detailed explanation of the key components and designs.
III-A Video Tokenization
Video tokenization involves encoding the raw video into visual tokens, which will be processed and then passed to the LLM for inference. We partition a long video into chunks, with each chunk represented by a set of compressed visual tokens. Such chunk tokens are stored in a memory to facilitate question-guided retrieval for efficient video question-answering.
@@ III-A Video Tokenization
We adopt the pretrained language-image model CLIP [20] to extract per frame visual token features. For the $j^{th}$ video chunk $V_{i}^{j}\in\mathbb{R}^{M\times H\times W\times 3}$, we obtain $M\times h\times w$ visual tokens $F_{i}^{j}\in\mathbb{R}^{M\times h\times w\times D}$ from the CLIP vision encoder, where $h=H/p,w=W/p$, $D$ denotes the number of channels. $p$ denotes the patch size ($p=14$ for ViT-L/14, $h=16$, $w=16$).
Chunk wise compression The preliminary token number of a chunk is large, which would result in a high computational burden and large memory requirement for the LLM even though we only select a few chunks as input to LLM. Therefore, we perform chunk-wise pooling to obtain reduced spatial and temporal tokens (see Fig. 3 for the illustration in Supplementary Section A).
Particularly, we found that reducing the spatial resolution by a factor of 4 leads to marginal difference in performance while this can significantly reduce the number of tokens by 75%. Therefore, we perform spatial average pooling with stride 2 to have $M$$\times$ $\bar{h}$$\times$ $\bar{w}$ tokens per chunk, where $\bar{h}$=$h/2$ and $\bar{w}$=$w/2$. This is equivalent to taking the CLIP features of reduced resolution as the extracted feature.
How to further reduce the number of tokens while preserving spatial-temporal features of a chunk? Motivated by Video-ChatGPT [1], we perform spatial-temporal pooling on the spatio-temporal features. Particularly, global spatial average pooling for each frame is performed and thus we obtain $M$ tokens for the $M$ frames (e.g., 4 tokens). Temporal pooling for each spatial position is performed to have $N=\bar{h}\times\bar{w}=8\times 8=64$ tokens. We have $N+M=64+4=68$ tokens $\bar{F}_{i}^{j}\in\mathbb{R}^{(N+M)\times D}$, with $D$ dimension for each token. Please see Supplementary for more details. Compared with the original 1024 in a chunk, the number of tokens has been reduced to 6.6%. We store the reduced chunk tokens in a memory to facilitate retrieval for the later LLM inference.
In contrast to Video-ChatGPT [1], which performs global spatial and temporal pooling over the entire video, ours with chunks is capable of preserving local details and is suited for long video reasoning.
III-B Question-guided Retrieval for Chunk Selection
@@
where the MLP block is a two-layer perceptron consisting of two fully connected layers and an ReLU activation function in between. This MLP transformation strengthens the correlation between the question representation and the potential corresponding chunk features for chunk selection. The MLP block is very lightweight with negligible number of parameters (i.e., 1.8M), while facilitating the question-aware retrieval for long video understanding. Note that achieving this is non-trivial, as we do not have any ground-truth locations of the question relevant chunks for supervision. The purpose of adding an MLP block to the text features is to enable learnable retrieval, i.e., the learnable selection of question-related video chunks. Another alternative design is to apply an MLP to the vision features during retrieval or apply MLPs to both. For a video question pair in inference, if we apply the MLP to the text encoder, we only need the inference through the MLP once. If we apply the MLP to the vision encoder, we need the inference through the MLP for every chunk. For simplicity, we only apply to the text feature.
To identify the question-related chunks, we measure the affinity between the question and the chunks. We define the similarity score between the question representation $\mathbf{q}_{i}$ and the $j^{th}$ chunk of video $V_{i}$ as the average cosine similarity between $\mathbf{q}_{i}$ and the $N+M$ chunk tokens as
$s_{i}^{j}=\frac{1}{N+M}\sum_{n=1}^{N+M}\frac{\mathbf{q}_{i}\cdot\mathbf{u}_{i,n}^{j}}{\lVert\mathbf{q}_{i}\rVert\lVert\mathbf{u}_{i,n}^{j}\rVert},$ (2)
@@
where $\mathbf{u}_{i,n}^{j}$ denotes the $n^{th}$ token in the $j^{th}$ chunk (of video $V_{i}$), $n=1,\ldots,N+M$. We rank the video chunks based on their similarity scores $s_{i}^{j}$ (where $j=1,\ldots,L_{i}$) and select the top $K$ most relevant chunks. The $K\times(N+M)$ visual tokens of these chunks are input to the LLM after a linear projection (a fully connected layer) on each token.
III-C End-to-End Optimization